OpenStudy (anonymous):

I'm new here and I need help with an additional mathematics course for GCSE's. My first question is in the first reply to this, regarding logarithms.

OpenStudy (anonymous):

Q. Using suitable substitutions, solve for x. \[2e ^{x} = 7\sqrt{e ^{x}} - 3\]

OpenStudy (anonymous):

e^x = t^2 you reduce your equation to 2t^2 = 7t -3

OpenStudy (anonymous):

now solve the quadratic eqn

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

@him1618 Yes! Thanks a bunch! :D